OpenStudy (anonymous):

There are 5 red and 9 green balls in a bag. If you randomly choose balls one at a time, with replacement, what is the probability of choosing 2 red balls and then 1 green ball?

OpenStudy (anonymous):

red= 1/7 green=1/14